# The Fraction of Clumpy Galaxies in JWST Surveys over $2<z<12$

### v1

@@ 5.3 Comparing Clumpy Galaxy Fractions between the HST and JWST Data
Clumps are identified in HST images using the algorithm by Sattari et al. (2023) and the detection criteria from Shibuya et al. (2016). These are the same algorithm and criteria as those used to search for clumps in the JWST images (Section 4.1). HST images have a pixel scale of 0.″06 pixel-1. We note that the criterion on the minimum number of contiguous pixels that must be spanned by a candidate clump depends on the HST instrument used: it is 5 for ACS images and 7 for WFC3 images. We search for clumps in HST data over $2<z<6.5$ to ensure that the coverage in rest-frame wavelength is as similar as possible between HST and JWST (see Section 3). Clumps are identified in HST bandpasses that span the rest-frame NUV at various redshifts. They are identified in the HST instruments/bandpasses as follows: ACS/F850LP over $2<z<3.5$; WFC3/F125W over $3.5<z<5$; and WFC3/F160W over $5<z<6.5$. Cutting on redshift removes 18 galaxies, with 1,227 remaining. We make sure that $f_{\rm{clumpy}}$ is measured using the same galaxies for both HST and JWST images. Therefore, any difference in $f_{\rm{clumpy}}$ is due to discrepancies in depth and resolution between HST and JWST images and not the selection of the sample of galaxies used to compare these data.
Figure 9: The fraction of the NUV luminosity per galaxy that originates in clumps. It is calculated only for clumpy galaxies and plotted as a function of stellar mass of the host galaxy in each panel. Clumps must be at least as bright as the lowest minimum fractional luminosity listed in Table 1 at a given redshift. Redshift increases from left to right and is indicated at the top of each panel. Measurements for individual galaxies are shown as small gray dots. At each redshift, medians are calculated in three bins in mass and are indicated with large black hexagons and horizontal error bars. The 16th and 84th percentiles of the fraction of NUV light in each bin are shown as the vertical error bars. Medians and percentiles are only calculated in mass bins that have at least 10 galaxies. The median fraction of NUV light from clumps is $\sim 15-20\%$ and changes little with redshift or the stellar mass of the host galaxy. Our measurements are consistent within uncertainties with those from Guo et al. (2015) over $2<z<3$, which are shown as black diamonds in the leftmost panel (see text for details).
In Figure 8, $f_{\rm{clumpy}}$ is plotted as a function of redshift for HST (indicated with dashed lines) and JWST data (indicated with solid lines). The limit on the luminosities of the clumps with respect to their hosts is varied in each panel. This is the minimum fraction of the host galaxy NUV luminosity a clump must have. It is highest in the left panel and selected to be 8% that of the host luminosity. This value is adopted from pre-JWST studies (Guo et al., 2015; Shibuya et al., 2016). It is lowered to 5% in the middle panel and 2% in the right panel. The latter is the lowest clump minimum fractional luminosity measured in this work (see Section 4.2).

### v2

@@ V.3 Comparing Clumpy Galaxy Fractions between the HST and JWST Data
Clumps are identified in HST images using the same algorithm and detection criteria described in Section IV.1, with one exception. Instead of using the segmentation map from JADES as the outer boundary in detecting clumps, as is done in Section IV.1, a limit is placed on the distance of a given clump to the center of its host to be less than $8\times R_{\textrm{half, F444W}}$, the half-light radius in F [^4] W, following Shibuya et al. (2016). For the purposes of the comparison here, this outer boundary criterion is also applied to JWST images. HST images have a pixel scale of 0.″06 pixel-1. We note that the criterion on the minimum number of contiguous pixels that must be spanned by a candidate clump depends on the HST instrument used: it is 5 for ACS and 7 for WFC3 images (Shibuya et al., 2016). We search for clumps in HST data over $2<z<6.5$ to ensure that the coverage in rest-frame wavelength is as similar as possible between HST and JWST (see Section III). Clumps are identified in HST bandpasses that span the rest-frame NUV at various redshifts with ACS/F850LP ($2<z<3.5$); WFC3/F125W ($3.5<z<5$); and WFC3/F160W ($5<z<6.5$). Cutting on redshift removes 18 galaxies, with 1,227 remaining. We make sure that $f_{\rm{clumpy}}$ is measured using the same galaxies for both HST and JWST images. Therefore, any difference in $f_{\rm{clumpy}}$ is due to discrepancies in depth and resolution between HST and JWST images and not the selection of the sample.
When measuring the brightness of clumps in the HST and JWST images, they are assumed to be point sources, instead of extended sources, as was found in Section IV.2. This is done because clumps were found to largely resemble point sources in works based on HST data (Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023). To treat clumps found in the JWST images on equal footing, they are also considered point sources only for the purpose of the comparison presented here. Details are given in Appendix B on the completeness of detecting clumps when treated as point sources and the aperture corrections needed.
JWST images are matched to the resolution of HST images using empirical PSFs in the HST bandpasses adopted from the 3D-HST survey (Skelton et al., 2014). They have a common pixel scale of 0.″06. They are upscaled by a factor of two to match the pixel scale of the JWST images. PSF-matching kernels are then created with pypher (Boucaud et al., 2016). After matching the resolution of HST and JWST, we performed photometry on the clumps in the same way as described in Section IV.2.
Aperture corrections for the HST images are 1.234, 1.337, 1.447 in GOODS-N and 1.225, 1.334, and 1.436 in GOODS-S for the F850LP, F125W, and F160W bandpasses, respectively. When matching JWST PSFs to those from HST, their aperture corrections agree to within 1%. When treating clumps as point sources in the JWST images, aperture corrections are 1.348, 1.307, and 1.299 in GOODS-N and 1.365, 1.328, and 1.330 in GOODS-S for the F090W, F115W, and F150W bandpasses, respectively.
Comparison of evolutionary trends for $f_{\rm{clumpy}}$ between HST and JWST is presented in Figure 9. The minimum fractional luminosity used to identify clumps is varied in each panel. It is highest in the left panel, corresponding to 8% of the host luminosity (Guo et al., 2015; Shibuya et al., 2016). It is lowered to 5% in the middle panel and 2% in the right panel. The latter is the lowest clump minimum fractional luminosity at which a complete sample of clumps can be detected in the JWST images, when treating them as point sources. $f_{\rm{clumpy}}$ measurements from JWST that are calculated with images at the original resolution are shown in the top row, while those at the resolution of HST are shown on the bottom.
Two conclusions may be drawn from Figure 9. First, at fixed minimum fractional luminosity, $f_{\rm{clumpy}}$ from JWST data changes little with resolution. Second, as the minimum fractional luminosity is progressively lowered, $f_{\rm{clumpy}}$ from HST barely changes, while that from JWST gradually changes.
To assess how much overlap there is in the clumps detected between the HST and JWST datasets, we count the number of clumps in the HST images with minimum fractional luminosity $\geq 8\%$. Next, for each clump detected with HST, we compare its $x$ and $y$ coordinates with those of clumps detected in JWST images with the same pixel scale as HST for the same host galaxy. We then count the number of clumps in the JWST images that lie within a given separation in pixels from its counterpart in the HST images. At a separation of 1.5 pixels, which corresponds to the PSF FWHM of F850LP (0.″09, Guo et al., 2013), 49% of clumps are matched between the two datasets. At a larger separation of 3 pixels, which is equal to the FWHM of the F160W PSF (0.″18), 72% of the clumps are matched. There are a few sources of uncertainty in these measurements. The center of the galaxy is determined separately for the HST and JWST data and may not be identical between the two. Detection images from JWST are significantly deeper than the HST, which could affect the determination of the centers of galaxies. Similarly, since a clump must possess sufficient contrast with the local background to be detected, differences in depth will affect how well clumps are detected and deblended. This is evident in our measurements. There are about 2.5 times more clumps in the convolved JWST images than in the HST (2,512 vs. 989) for the same galaxies and at a minimum fractional luminosity of $\geq 8\%$.
We attribute the results discussed above mostly to the difference in sensitivity between HST and JWST. The latter can detect point sources that are $\sim 2-5$ times fainter than the former, depending on the wavelength [^1] https://jwst-docs.stsci.edu/jwst-opportunities-and-policies/jwst-general-science-policies/guidelines-for-proposals-where-jwst-and-hst-overlap-in-capabilities#gsc.tab=0. It is therefore not surprising that $f_{\rm{clumpy}}$ is systematically higher when measured with JWST, when the faintest clumps are considered. Furthermore, our finding that resolution has a secondary ($\leq$9 %) effect on $f_{\rm{clumpy}}$ is similar to what Guo et al. (2012) found with HST images for a sample of 40 bright clumps in star-forming galaxies over $1.5<z<2.5$ and with stellar masses $M_{\star}\gtrsim 10^{10}M_{\odot}$.
VI Clumpy Fraction as a Function of Stellar Mass
The fraction of clumpy galaxies as a function of stellar mass of their host galaxy is plotted in redshift bins in Figure 10, left panel. The fraction is calculated in three bins, spanning stellar masses $8<\log\left(M_{\star}/M_{\odot}\right)<9$, $9<\log\left(M_{\star}/M_{\odot}\right)<10$, and $10<\log\left(M_{\star}/M_{\odot}\right)<11$. Fractions are only computed for bins that have at least 10 galaxies.
At all redshifts, $f_{\rm{clumpy}}$ increases with the stellar mass. At a fixed mass, as redshift increases, $f_{\rm{clumpy}}$ decreases. At the lowest redshifts, $z\sim 2.75$, $f_{\rm{clumpy}}$ increases from $\sim 20\%$} at $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$ to $\sim 70\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$ and $\sim 85\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 10.5$. At higher redshifts, $f_{\rm{clumpy}}$ increases with stellar mass at a rate similar to that at $z\sim 2.75$. At $z\sim 4.75$, it grows from $\sim 5\%$ at the lowest masses to $\sim 70\%$ at the highest masses. At $z\sim 5.75$, it rises from $\sim 1\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$ to $\sim 20\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$. At $z\sim 6.75$, a complete sample of clumps is only detected at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$, for which $f_{\rm{clumpy}}$ is $\sim 17\%$.
The above results were found by selecting clumps using the minimum fractional luminosities listed in Table 1, which vary with stellar mass and redshift. However, the correlation between stellar mass and $f_{\rm{clumpy}}$ still holds when selecting clumps with the same minimum fractional luminosity at all masses. The fraction of clumpy galaxies is calculated again for galaxies with stellar mass $\log\left(M_{\star}/M_{\odot}\right)\sim 8.5$, with the minimum fractional luminosities this time being the same as those applied to higher-mass galaxies. The result is shown in the left panel in Figure 10. The effect of lowering the minimum fractional luminosity for the lowest-mass galaxies is an increase in $f_{\rm{clumpy}}$ by $\sim 10-20\%$ at fixed redshift.
In Figure 10, right panel, we compare the correlation between stellar mass and $f_{\rm{clumpy}}$ at $z\sim 2.75$ with that measured by other studies over $2<z<3$. These measurements are shown as various symbols. Guo et al. (2015) and Sattari et al. (2023) found that, as stellar mass increases, $f_{\rm{clumpy}}$ decreases. It was found to be $\sim 60\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$ and $\sim 40\%$ at $\log\left(M_{\star}/M_{\odot}\right)\sim 11$. Both studies used HST data covering the rest-frame UV and developed detection algorithms to search for clumps. In contrast, Huertas-Company et al. (2020) found that it has a positive slope, with $f_{\rm{clumpy}}$ increasing from $\sim 20\%$ to $\sim 30\%$ with as stellar mass increases from $\log\left(M_{\star}/M_{\odot}\right)\sim 9$ to 11.
Results from the present work at $z\sim 2.75$ disagree with those from other studies at similar redshifts, finding a positive correlation between stellar mass and $f_{\rm{clumpy}}$.
VII Comparison with Hydrodynamic Simulations
Here we discuss the implication of our results to galaxy formation scenarios. There are two key results. First, the fraction of clumpy galaxies, $f_{\rm{clumpy}}$, increases with decreasing redshift over 2<z<9{2<z<9}. For galaxies with stellar mass $M{\star}\geq 10^{9}M_{\odot}$, $f_{\rm{clumpy}}$ is $\sim 10\%$} at $z\sim 7.75$} and $\sim 70\%$} at $z\sim 2.75$. We find relatively higher $f_{\rm{clumpy}}$ in JWST-selected samples compared to those selected by HST. This is due to the higher sensitivity of the JWST observations (Section V and Figure 8). Second, at a fixed redshift, as stellar mass increases, the fraction of clumpy galaxies increases too (Section VI and Figure 10).
In this Section, we compare our results with simulations of clumpy galaxies. This is done to identify potential physical mechanisms for their formation. We compare with zoom-in hydrodynamic cosmological simulations from Mandelker et al. (2017) and Nakazato et al. (2024). The former studied 34 galaxies over $1\lesssim z\lesssim 6$ from the VELA simulations (Ceverino et al., 2014; Zolotov et al., 2015), which had stellar masses of $\log\left(M_{\star}/M_{\odot}\right)\gtrsim 9$ at $z=2$, with clumps largely forming through VDI. The latter analyzed 62 galaxies over $5.5<z<9.5$ from the FirstLight simulations (Ceverino et al., 2017), which had stellar masses in the range $(0.5-6)\times 10^{10}M_{\odot}$ at $z=5$. Most clumps in these simulations formed from major mergers. The $f_{\rm{clumpy}}$  estimates and predictions are compared in Figure 11.
Figure 11: Observed $f_{\rm{clumpy}}$ values as a function of redshift in three stellar mass bins compared with predictions from simulations. Uncertainties are shown as shaded regions. The fraction of clumpy galaxies that formed from violent disk instabilities (VDI; Mandelker et al., 2017) at $z<6$ is shown as gray, red, and green shaded regions, in the order of increasing stellar mass. Predictions for clumpy galaxies that formed from major mergers over $6\lesssim z\lesssim 9$ (Nakazato et al., 2024) are shown as a purple line and shaded region. Observed pair fractions as a function of stellar mass and redshift are shown as thin gray lines (Puskás et al., 2025). The results from this work are consistent with $f_{\rm{clumpy}}$ from simulations based on VDI and those based on major mergers at redshifts and masses where they overlap. The observed pair fractions at $\log\left(M_{\star}/M_{\odot}\right)\leq 9$ are also consistent with the observed $f_{\rm{clumpy}}$ measured here at similar masses and redshifts.
The $f_{\rm{clumpy}}$ measurements are consistent with predictions from Mandelker et al. (2017) at $z\lesssim 5.75$ and $\log\left(M_{\star}/M_{\odot}\right)\geq 9$. The predictions increase rapidly with redshift from $\sim 20\%$ at $z\sim 1$ to $\sim 50\%$ at $z\sim 2$, above which they asymptote. They have uncertainties of $\sim 20\%$, which result from Poisson noise due to a limited number of galaxies examined. While our $f_{\rm{clumpy}}$ measurements peak at $\sim 70-80\%$} at $z\sim 2.75$ and decline with increasing redshift, the model predictions do not change with redshift, when taking into account the uncertainties. This discrepancy is likely due to differences in the way clumps are identified. Mandelker et al. (2017) searched for clumps in projected density maps of their galaxies and computed $f_{\rm{clumpy}}$ using those with stellar masses $10^{7}~M_{\odot}$ and SFRs of at least 0.1% those of their host galaxies’ disks. In this work, clumps are detected in rest-frame NUV images and selected to be compact objects that are at least as bright as a fraction of their hosts’ NUV luminosity, depending on redshift and stellar mass of the host galaxy. In both studies, clumps are selected to be off-center.
Our $f_{\rm{clumpy}}$ measurements are also consistent with those from Nakazato et al. (2024) at $z\sim 7.75$} and $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$}. At these redshifts and stellar masses, we find $f_{\rm{clumpy}}$ to be 1313%. A similar, near-constant value of $\sim 10\%$ is obtained by Nakazato et al. (2024) over $6<z<9$. Galaxies studied in that work had stellar masses of $\lesssim 10^{9}~M_{\odot}$ at $z\geq 8.5$. Therefore, trends in $f_{\rm{clumpy}}$ at the intermediate masses considered in this work are consistent with those from Nakazato et al. (2024) at redshifts where they overlap. However, clumps were identified differently between these two works. In Nakazato et al. (2024), clumps were found in stellar mass surface density maps of young (<10<10 Myr) stars. Detectable clumps had stellar masses of $\gtrsim 10^{7}~M_{\odot}$. This is about an order of magnitude lower than the faintest clumps that can be detected in this work at $z\sim 7.75$} and $\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$, assuming clumps have similar NUV mass-to-light ratios as their hosts at these redshifts and stellar masses. Furthermore, in Nakazato et al. (2024), galaxies were defined to be clumpy if they had at least two clumps.
It is unclear from the simulations which processes dominate the formation of clumpy galaxies at other combinations of redshift and stellar mass, particularly at the lowest masses. Predictions for $f_{\rm{clumpy}}$ for galaxies at these masses and redshifts are needed. Predictions from simulations that produce clumps through VDI are needed out to higher redshifts in order to distinguish clumps formed through VDI and those from mergers. Our $f_{\rm{clumpy}}$ measurements at intermediate stellar masses ($9<\log\left(M_{\star}/M_{\odot}\right)<10$) are about two times higher than those from Nakazato et al. (2024) at $z\sim 5.75$. This suggests that major mergers may not significantly contribute to clumpy galaxy formation at these masses and redshifts (see below for support for this from observations). However, confirming this will require $f_{\rm{clumpy}}$ measurements from simulations of clumpy galaxies that form via other mechanisms at these redshifts.
VIII Discussion
In this section, we discuss observations of the morphologies, gas kinematics, and pair fractions of star-forming galaxies with similar redshifts and stellar masses as those studied in this work. This is done to assess the plausibility of the mechanisms considered in the previous section, wherein our measurements of $f_{\rm{clumpy}}$ were compared with predictions from simulations. The comparison suggests that VDI and mergers may be the dominant mechanisms for clumpy galaxy formation, depending on redshift and stellar mass, in agreement with other studies (e.g., Guo et al., 2015; Shibuya et al., 2016). The dominant mechanism for forming clumps appears to be VDI at $z\lesssim 5.75$ and at intermediate-to-high stellar masses ($\log\left(M_{\star}/M_{\odot}\right)\geq 9$) and mergers at $z\gtrsim 7.75$} and intermediate stellar masses ($\log\left(M_{\star}/M_{\odot}\right)\sim 9.5$}). We find that the observations listed above largely support the scenario described here.
For clumps to form through VDI, their host galaxies must possess marginally stable disk components (e.g., Bournaud et al., 2007; Agertz et al., 2009; Dekel et al., 2009b; Ceverino et al., 2010; Mandelker et al., 2014). Star-forming galaxies over $2<z\lesssim 6$ and $\log\left(M_{\star}/M_{\odot}\right)>8$ have typical Sérsic indices of $n\simeq 1$ (van der Wel et al., 2014; Shibuya et al., 2015, 2016; Nedkova et al., 2021; Ormerod et al., 2024; Sun et al., 2024; Varadaraj et al., 2024; Ward et al., 2024; Miller et al., 2025; Yang et al., 2025), which suggests they have disk-like structures. At higher redshifts, Sérsic indices are generally larger, which implies disk-like structure becomes less common (Sun et al. 2024; Miller et al. 2025; see also Ferreira et al. 2023; Kartaltepe et al. 2023).